OpenStudy (anonymous):

What is the value of the expression? 5^-6 * 5^2/5^8

OpenStudy (zale101):

Here are some basic rules that you need to consider \(\Large a^b*a^c=a^{b+c}\) \(\LARGE \frac{ a^b}{a^c}=a^{b-c}\)

OpenStudy (zale101):

First, work on the numerator by using \(\Large a^b*b^c=a^{b+c}\) Then do the division exponent rule.

OpenStudy (zale101):

Makes sense?

OpenStudy (anonymous):

Hmmm kinda

OpenStudy (zale101):

Try working this out on here or on paper then type down what you got.

OpenStudy (anonymous):

Otay :)

OpenStudy (zale101):

\(\LARGE \frac{5^{-6} * 5^2}{5^8}\)=?

OpenStudy (zale101):

\(\LARGE \frac{5^{-6+2}}{5^8}\)

OpenStudy (zale101):

And i'll leave the rest to you

OpenStudy (anonymous):

Okay thank you :)

OpenStudy (zale101):

Anytime~

OpenStudy (anonymous):

So is it 5^-12

OpenStudy (anonymous):

I forgot to add the negative in 5^8 it was actually 5^-8

OpenStudy (anonymous):

hello? ;_;

OpenStudy (zale101):

\(\LARGE \frac{5^{-6+2}}{5^{-8}}=\frac{5^{-4}}{5^{-8}}=5^{-4-(-8)}=5^{-4+8}=5^{4}\)

OpenStudy (anonymous):

Oh thank I understand now :) ♥♥♥

OpenStudy (zale101):

No problem :)